OpenStudy (user917):

Is lim (h -> 0) of h*f(x) = 0 always true?

OpenStudy (anonymous):

where is the function?

OpenStudy (user917):

function is R to R and differentiable that the only thing I know

OpenStudy (user917):

but isn't 0 * anything = 0 ? @magepker728

OpenStudy (anonymous):

what is hf(x)?

OpenStudy (anonymous):

or u mean f(h)?

OpenStudy (user917):

@magepker728 h*f(x)

OpenStudy (welshfella):

i think its true.

OpenStudy (bobo-i-bo):

Assuming f nor x are functions of h, then it has to be true, since the limit would then treat f(x) as a constant. But this is not necessarily the case if either f or x is a function of h. For example: Assuming \(x \neq 0\), if \(f(x)= \frac x h\) or if \(f(x)=x\) and \(x=\frac 1 h\), then \(\lim(hf(x))=\lim(\frac {hx} h)=\lim x=x \neq 0\)